### v1

@@ 6.3 Non-linearity
In Figure 8, we plot the effect on GVA (y-axis) against the amount received from the Cohesion Fund as a share of GVA for each treated region, i.e. the treatment intensity. Each black dot represents the average “treatment effect” for each treated region. A spline is fitted to these dots represented by the blue line. Interestingly, a nonlinear relationship between the impact on GVA and the cohesion funds received as a share of GVA is revealed. This echoes the concept of a maximum desirable level of transfers of Becker et al. (2012), based on which beyond a certain level of treatment intensity the impact declines.
Figure 8: Nonlinear effect of Cohesion Funds
[…]
At relatively low funding as a share of GVA, the effect on GVA appears small, negligible or even negative. However, as funding increases, the effect grows, reaching its peak when the Cohesion Fund accounts for 0.58% of the region’s GVA over the first seven years. Beyond this point, we observe diminishing returns, where the effect of cohesion funds gradually declines and eventually becomes zero or even negative at higher funding shares of GVA. Similarly, the treatment effect peaks when the CF share reaches 0.94% of a region’s GVA for the fourteen year period On average the maximum effect is achieved when the CF share is equal to 0.54% of a region’s GVA.
This nonlinear relationship suggests that the EU policymaker should allocate this type of funding with caution. On the one hand, increased funding can generate positive effects, but only up to a certain upper limit where the impact reaches its peak. Beyond this point, additional funding leads to diminishing returns and may even become counterproductive above a certain threshold. Our analysis enables us to quantify both the upper bound and the threshold beyond which further funding results in negative effects.
@@ 7 Conclusions
This paper studies the impact of Cohesion Fund payments on regional economic performance, focusing on the heterogeneous responses of the EU’s regions to this sub-national transfers programme. In particular, relying on the matrix completion technique, it disentangles both the region- and the time-specific effects of the policy. Our main results are that these transfers exert, on average, a positive and persistent effect on both output and investment, with the majority of the gains materializing in the first seven years under treatment. Moreover, the relatively poorer regions are the ones reaping the benefits of the programme. Lastly, we uncover a nonlinear relationship between funding intensity and the size of the effect.
@@
Appendix A Technical details of empirical strategy
@@
Here, we provide definitions for terms used in Section 5. A real-valued random variable $epsilon$ is $\sigma$-sub-Gaussian if for all $s\in\mathbb{R}$,
@@
The singular value decomposition (SVD) of a $N\times T$ matrix A writes $A=S\Sigma R^{\top}$, where $S$ and $U$ are $N\times N$ and $T\times T$ matrices respectively and $\Sigma$ is $N\times T$ rectangular diagonal matrix with elements denoted by $\sigma_{i}(A)$, $i=1,\ldots,\min(N,T)$. The Frobenius norm of a matrix A is defined as
@@
We assume $K=2$ latent factors in (3) and $\lambda_{i1}=1$, $\lambda_{i2}\in\mathbb{R}$, $f_{1t}\in\mathbb{R}$ and $f_{2t}\in\mathbb{R}$ as well as $\tau_{it}=\tau$ for each $i=1,\ldots,N$ and $t=1,\ldots,T$ then (3) becomes
$y_{it}=\tau D_{it}+f_{1t}+\lambda_{2i}+\epsilon_{it},$
@@
Estimation of the low rank matrix $L$
Here, we describe briefly the methodology for solving the minimization problem defined by equation (6) in order to estimate the matrix $L$; see Athey et al. (2021) for more details.
First we note an estimator of $L$ could be also calculated by minimizing the sum of squared differences,
$\min_{\mathbf{L}}\frac{1}{|\mathcal{O}|}\sum_{(i,t)\in\mathcal{O}}(Y_{it}-L_{it})^{2}=\min_{\mathbf{L}}\frac{1}{|\mathcal{O}|}\|\mathbf{P}_{\mathcal{O}}(Y-L)\|_{F}^{2},$ (8)
where $\mathcal{O}\subset\{1,\dots,N\}\times\{1,\dots,T\}$ is the set of pairs $(i,t)$ such that $y_{it}$ is an observed entry in the matrix $Y$ with potential untreated outcomes. However, the objective in (8) is not useful, as the estimator simply sets $L_{it}=y_{it}$ for observed entries. To address this, Athey and Imbens (2017) introduce the regularization term $\lambda\|L\|$ in (8) and even more to avoid the regularization of fixed effects the modify the minimization in (6) to obtain an estimator of the form
$(\hat{L},\hat{\Gamma},\hat{\Delta})=\arg\min_{L,\Gamma,\Delta}\left\{\frac{1}{|\mathcal{O}|}\|P_{\mathcal{O}}(Y-L-\Gamma\mathbf{1}_{T}^{\top}-\mathbf{1}_{N}\Delta^{\top})\|_{F}^{2}+\lambda\|L\|_{*}\right\},$ (9)
where $\mathbf{1}_{N}$ and $\mathbf{1}_{T}$ are $N$- and $T$- dimensional vector with ones.
Importantly, the resulting estimator can be computed using fast convex optimization programs, such as the soft-impute algorithm (Mazumder et al., 2010) which is briefly described as follows. For any given matrix $A$ with SVD $A=S\Sigma R^{\top}$ Athey et al. (2021) define the matrix shrinkage operator
$\operatorname{shrink}_{\lambda}(\mathbf{A})=S\tilde{\Sigma}R^{\top},$ (10)
where in $\tilde{\Sigma}$ is equal to $\Sigma$ but each singular value $\sigma_{i}(A)$ of the latter is replaced by $\max(\sigma_{i}(A)-\lambda,0)$. Using this, we initialize
$L_{1}(\lambda,\mathcal{O})=\mathbf{P}_{\mathcal{O}}(\mathbf{Y}),$
and iteratively update
$L_{k+1}(\lambda,\mathcal{O})=\operatorname{shrink}_{\frac{\lambda|\mathcal{O}|}{2}}\left\{P_{\mathcal{O}}(Y)+P_{\mathcal{O}}^{\perp}(L_{k}(\lambda,\mathcal{O}))\right\}.$ (11)
This iteration continues until convergence. Finally, the optimal $\lambda$ is selected via cross-validation where $\mathcal{O}$ is randomly splitted into $K$ subsets, ensuring that each subset preserves the observed fraction of data. A sequence of candidate values $\lambda_{1}>\lambda_{2}>\dots>\lambda_{L}=0$ is tested, and the value minimizing the average squared error on hold-out data is chosen. To expedite computation, each $\hat{L}(\lambda_{i},\mathcal{O}_{k})$ can serve as a warm-start for computing $\hat{L}(\lambda_{i+1},\mathcal{O}_{k})$.
Appendix B Data
@@
As can be gleaned from Figure B.1, the correlation between the modeled and the actual payments, as shares of GVA, is –on average– very high, with a correlation coefficient equal to 0.93 (see, also, Fiuratti et al. (2024) for additional analysis).
Figure B.1: Correlation between modelled and raw Cohesion Fund payments
@@
Appendix C Tables and figures
C.1 Observed and counterfactual GVA levels and growth rates
Figure C.1: Observed and estimated counterfactual average GVA in logs.
Figure C.2: Observed and estimated counterfactual average GVA growth rates.
C.2 Tables of growth rates excluding outliers
Table C.1: Log-difference of GVA per capita by programming period (%) excluding outliers
Period Regions Observed Counterfactual
1994-1999 32 13.84 18.01
2000-2006 87 22.2 13.8
2007-2013 87 -1.66 0.69
2014-2020 71 22.4 15.11
  • Notes: The table depicts the difference between the end and beginning of period log of GVA per capita. As such, it corresponds to the overall period growth rate (not the annual average).
  • The 1994-1999 period actually corresponds to the 1993-1999 calendar year, as all the regions received a small proportion of the Cohesion Fund envelope in 1993.
  • The decline in regions in the 2014-2020 period is due to Spain not being eligible for the Cohesion Fund.
Table C.2: Log-difference of GVA per capita, years under treatment (%) excluding outliers
Years Regions Observed Counterfactual
7 87 20.9 14.2
14 87 9.87 9.33
21 83 4.31 1.78
  • Notes: For additional notes, see Table
Table C.3: GVA growth rates by years under treatment
Percentile Initial GVA Years under treatment $g_{y}^{\textrm{treated}}$, in % $g_{y}^{\textrm{counter}}$, in %
10th 4137.21 7 Years 25.7 -9.03
25th 5273.65 23.22 9.33
50th 8416.783 18.38 9.54
75th 11928.78 8.06 13.27
90th 17153.59 13.29 22.15
10th 4137.21 14 Years -22.33 1.79
25th 5273.65 17.14 8.98
50th 8416.783 -6.01 1.47
75th 11928.78 20.04 11.52
90th 17153.59 11.66 15.88
  • Notes: The last two columns depict the overall period growth rate for the cases in which the region is treated and the counterfactual, respectively.
    Initial GVA refers to the GVA per capita in the first year of the treatment

### v2

@@ 6.4 Non-linearity
Figure 13 plots the estimated Average Treatment Effect on the Treated ($\widehat{ATT}$) against the intensity of Cohesion Fund transfers –measured as the share of real Cohesion Fund expenditures in real GVA– for treated regions. Each curve corresponds to different time horizons, i.e., 1 year before treatment as well as 7, 14, and 21 years post-treatment. The relationship between treatment intensity and economic impact is clearly non-linear across all time horizons.
As expected, one year prior to the treatment, the estimated $\widehat{ATT}$ remains close to zero across all levels of funding intensity. As the post-treatment horizon extends, however, the non-linear relationship between treatment intensity and the estimated effect $\widehat{ATT}$ becomes increasingly pronounced (see the changing curvature of the lines as the post-treatment year varies from 1 to 21). Over time, the treatment intensity, i.e., the CF as a share of GVA, which generate the maximum estimated $\widehat{ATT}$– depicted by the dots on each curve– rises. Specifically, in year 7, the optimal treatment intensity is approximately 0.6% of GVA, at which point the estimated effect reaches 25%. Beyond this threshold, additional funding produces diminishing economic returns. In subsequent periods, both the optimal intensity and the peak effect rise; however, once funding exceeds the respective threshold, further transfers lead to a sharp decline in the estimated effect. The optimal treatment intensity is estimated at 0.7% and 0.86% for post-treatment years 14 and 21, respectively, yielding maximum $\widehat{ATT}$ effects of 0.32 and 0.47 in logs, or 35% and 60.4%, respectively.
These results support the hypothesis of a maximum desirable level of transfers proposed by Becker et al. (2012), suggesting that while moderate levels of EU funding can stimulate growth, when transfers exceed a threshold, they can become inefficient or even counterproductive. Our results extend their analysis by emphasizing the time-varying dimension. [^23] Our results are in line with Becker et al. (2012) and Cerqua and Pellegrini (2018) who estimate that the optimal transfer intensity at 0.4% and the maximum desirable level at 1.3% Specifically, Figure 13 quantifies how the optimal level of treatment intensity –measured as funds relative to GVA– evolves as regions progress from the initial inclusion in the program to relatively more mature stages of the program implementation. Also, the right downward-sloping segment of the curves also provides an indication of how rapidly these effects decay as funding intensity increases.
This analysis offers quantitative guidance for policy. That is, funding allocations should be scaled over time to remain within an effective range, avoiding the diminishing or even adverse effects observed at higher treatment intensities.
This observed shift in the treatment intensity threshold and the resulting maximum $\widehat{ATT}$ effect is likely correlated with two factors: (i) the absorption capacity of the regions and (ii) the irregular disbursements of the funds. Overall, these results highlight the value added of the methodology for policy evaluation adopted in this paper, which allows for the calculation of time-specific effects. Previous research, see Becker et al. (2012), Cerqua and Pellegrini (2018), provided clear evidence of a non-linear relationship for an entire programming period or combinations of periods –often constrained by limited data availability– and therefore could not capture the temporal shift in the maximum desirable level of transfers.
Figure 13: Non-linear effect of the Cohesion Fund
@@ 7 Conclusions
This paper studies the impact of Cohesion Fund payments on regional economic performance, focusing on the heterogeneous responses of the EU’s regions to this sub-national transfers programme. In particular, relying on the matrix completion technique, it disentangles both the region- and the time-specific effects of the policy. Our main results are that these transfers exert, on average, a positive and persistent effect on both output and investment, with the majority of the gains materializing in the first seven years under treatment. Moreover, the relatively poorer regions are the ones reaping the benefits of the programme. Lastly, we uncover a non-linear relationship between funding intensity and the size of the effect.
@@
Appendix A Review of the literature
Table B provides a schematic review of the empirical literature concerned with the ex-post evaluation of cohesion policy transfers. We note that this section simply aims at providing an introduction to this voluminous literature by discussing some representative contributions; by no means should it be considered as being a thorough literature review.
The main structural feature of this selective review will be the methodological approach employed and, secondarily, the type of data used for the cohesion transfers. More specifically, the focus in on whether a form of a policy evaluation technique (e.g. a Regression Discontinuity Design) or more traditional (panel) econometrics approach was considered.
The top panel of Table B collects those contribution that focused on a causal inference approach. As this part was elaborated on in the main body of the paper, we briefly mention here that in terms of the empirics either an RDD or a dose response model was employed, whereas in terms of cohesion data either a binary indicator related to region eligibility or data on transfer commitments and transfer payments provided to the authors by the European Commission - Directorate General for Regional and Urban Policy (DG REGIO).
As already mentioned in the main text, apart from the research stream focused on applying causal inference techniques for the ex-post evaluation of Cohesion policy, more traditional econometric approaches have also been utilized.
The bulk of this literature usually adopts a panel data framework to assess the effects of cohesion policy on regional economic performance. This body of work typically exploits variation over time and across regions to estimate the long-run impacts of EU transfers. Within this strand, some studies rely on instrumental variables or spatial econometric techniques to address endogeneity concerns and account for spatial interdependencies (see, e.g., Di Caro and Fratesi 2022, Mohl and Hagen 2010). Others incorporate institutional and governance dimensions into their empirical designs, highlighting how these factors condition policy effectiveness (e.g. Rodríguez-Pose and Garcilazo 2015).
More recent contributions shift the focus to short-run multiplier effects of cohesion spending, estimating how transfers influence regional output dynamics in the immediate aftermath of disbursement (e.g. Canova and Pappa 2025).
Policy evaluation for EU funding impact assessment
RDD
Becker et al. (2013) 186-251 NUTS2(p) 1989-2006 Binary GDP pc growth Fuzzy RDD Homogeneous (HLATE)
Becker et al. (2018) 259 NUTS2 1989-2013 Binary GDP pc growth Fuzzy RDD Homogeneous
Crescenzi and Giua (2020) NUTS2 regions in 4 countries 2000-2014 Binary GDP pc growth Spatial RDD Heterogeneous (partially)
Gagliardi and Percoco (2017) 257 NUTS2, 1233 NUTS3 2000-2006 Binary GDP pc growth Fuzzy RDD Homogeneous
Percoco (2017) 257 NUTS2, 1233 NUTS3 2000-2006 Binary GDP pc growth Fuzzy RDD Homogeneous(HLATE)
Dose-response
Mohl and Hagen (2008) 122 NUTS1/NUTS2 1995-2005 Payments GDP pc growth Dose-response (GPS) Homogeneous
Panel data
Long-run analysis
Multiplier analysis
Canova and Pappa (2025) 281 NUTS2 1990-2018 Cohesion database GVA pc growth Local Projections Homogeneous panel
Coelho (2019) 238 NUTS2 2000-2013 Payments, Commitments GDP pc growth Panel IV Homogeneous panel
Durand and Espinoza (2021) EU27 + UK 1994-2018 Cohesion database GDP pc growth Local Projections Homogeneous panel
Fiuratti et al. (2024) NUTS2 of 25 EU countries 1994-2018 Cohesion database GDP pc growth Local Projections Homogeneous panel
  • Homogeneity: Partial means that the paper includes results based on groupings of regions. HLATE denotes the case in which a heterogeneous LATE is estimate in an RDD setup, by allowing the treatment effect to vary along covaraites that do not influence the treatment status.
  • Weighting matrix: In order to preserve the readability of the table, the various weighting schemes employed are mentioned here: k-nearest neighbor, Gaussian distance decay function.
Appendix B Technical details of empirical strategy
@@
Here, we provide definitions for terms used in Section 5. A real-valued random variable $\epsilon$ is $\sigma$-sub-Gaussian if for all $s\in\mathbb{R}$,
@@
The singular value decomposition (SVD) of a $N\times T$ matrix $A$ writes $A=S\Sigma R^{\top}$, where $S$ and $R$ are $N\times N$ and $T\times T$ orthonormal matrices, respectively, and $\Sigma$ is an $N\times T$ rectangular diagonal matrix with singular values $\sigma_{i}(A)$, $i=1,\ldots,\min(N,T)$.
The Frobenius norm of a matrix $A$ is defined as
@@
We assume $K=2$ latent factors in (3) and $\lambda_{i1}=1$, $\lambda_{i2}\in\mathbb{R}$, $f_{1t}\in\mathbb{R}$ and $f_{2t}\in\mathbb{R}$ as well as $\tau_{it}=\tau$ for each $i=1,\ldots,N$ and $t=1,\ldots,T$, then (3) becomes
$y_{it}=\tau D_{it}+f_{1t}+\lambda_{2i}+\epsilon_{it},$ (7)
@@
Estimation of the low-rank matrix $L$ and $\beta$
The minimization in (LABEL:eq:min_with_covariates) is a convex problem and can be solved iteratively. A popular approach, as discussed in Athey et al. (2021), is to alternate between the estimation of $\beta$ and the low-rank matrix $L$. That is:
  1. 1.
    Initialize $\beta^{(0)}$, and compute residual matrix $R^{(0)}=Y-\mathbb{X}(\beta^{(0)})$.
  2. 2.
    Solve
    $(\hat{L}^{(0)},\hat{\Gamma}^{(0)},\hat{\Delta}^{(0)})=\arg\min_{L,\Gamma,\Delta}\left\{\frac{1}{|\mathcal{O}|}\|P_{\mathcal{O}}(R^{(0)}-L-\Gamma\mathbf{1}_{T}^{\top}-\mathbf{1}_{N}\Delta^{\top})\|_{F}^{2}+\lambda\|L\|_{*}\right\}.$
  3. 3.
    Update $\beta^{(1)}$ by solving the weighted least squares problem on the residualized data $Y-\hat{L}^{(0)}-\hat{\Gamma}^{(0)}\mathbf{1}_{T}^{\top}-\mathbf{1}_{N}(\hat{\Delta}^{(0)})^{\top}$.
  4. 4.
    Repeat until convergence.
Each update of $L$ can be performed via soft-thresholding of singular values using the matrix shrinkage operator as described below.
Matrix shrinkage and iterative soft-thresholding
For any matrix $A$ with SVD $A=S\Sigma R^{\top}$, define the shrinkage operator
$\operatorname{shrink}_{\lambda}(\mathbf{A})=S\tilde{\Sigma}R^{\top},$
where $\tilde{\Sigma}$ replaces each singular value $\sigma_{i}(A)$ with $\max(\sigma_{i}(A)-\lambda,0)$.
Using this, the soft-impute algorithm begins with
$L_{1}=P_{\mathcal{O}}(Y-\mathbb{X}(\hat{\beta})-\hat{\Gamma}\mathbf{1}_{T}^{\top}-\mathbf{1}_{N}\hat{\Delta}^{\top}),$
and updates iteratively:
$L_{k+1}=\operatorname{shrink}_{\frac{\lambda|\mathcal{O}|}{2}}\left\{P_{\mathcal{O}}(Y-\mathbb{X}(\hat{\beta})-\hat{\Gamma}\mathbf{1}_{T}^{\top}-\mathbf{1}_{N}\hat{\Delta}^{\top})+P_{\mathcal{O}}^{\perp}(L_{k})\right\}.$
Tuning parameter selection
The regularization parameter $\lambda$ is selected via cross-validation. The set of observed entries $\mathcal{O}$ is randomly split into $K$ folds, and a grid of candidate values $\lambda_{1}>\lambda_{2}>\cdots>\lambda_{L}=0$ is tested. The average squared prediction error on hold-out folds is computed, and the value minimizing it is selected. To improve convergence, warm-starts are used: the solution for $\lambda_{i+1}$ is initialized using the solution from $\lambda_{i}$.
Appendix C Data and comparisons
@@
As can be gleaned from Figure C.1, the correlation between the modeled and the actual payments, as shares of GVA, is –on average– very high, with a correlation coefficient equal to 0.93 (see, also, Fiuratti et al. (2024) for additional analysis).
Figure C.1: Correlation between modelled and raw Cohesion Fund payments
@@
Comparison with TWFE
Figure C.2 compares the estimated Average Treatment Effects (ATT) over time using two different methodologies; the traditional Two-Way Fixed Effects (TWFE) estimator and the Matrix Completion with Nuclear Norm Minimization (MC-NNM) method proposed by Athey et al. (2021). Both estimators produce qualitatively similar trajectories. The overall similarity in results lends credibility to the findings and suggests a degree of robustness in the estimated effects of the CF program.
That said, the matrix completion estimates tend to be slightly higher than those from TWFE, especially in the medium and long run. This difference is consistent with the fact that TWFE can bias estimates downward in the presence of treatment effect heterogeneity or staggered treatment adoption—both of which are relevant in this context. A key advantage of the matrix completion approach is that it does not impose constant treatment effects and instead accommodates complex patterns of response across units and over time. Importantly, the matrix completion framework also enables the heterogeneity analysis discussed in earlier figures.
Figure C.2: $\widehat{ATT}_{t}$ estimated by using the two way fixed effects (TWFE) against $\widehat{ATT}_{t}$ presented by Figures 4 and 4 for log-GVA per capita (left) as well as for log-gross fixed capital formation (right).
Appendix D Tables and figures
D.1 Observed and counterfactual GVA levels and growth rates
Figure D.1: Observed and estimated counterfactual average GVA in logs.
Figure D.2: Observed and estimated counterfactual average GVA growth rates.
D.2 Treatment Status
Figure D.3: Treatment Status by Region and Year